# Galley: Modern Query Optimization for Sparse Tensor Programs

### v4

@@ Abstract.
The tensor programming abstraction has become a foundational paradigm for modern computing. This framework allows users to write high performance programs for bulk computation via a high-level imperative interface. Recent work has extended this paradigm to sparse tensors (i.e., tensors where most entries are not explicitly represented) with the use of sparse tensor compilers. With these systems, users define the semantics of the program and the algorithmic decisions in a concise language that can be compiled to efficient low-level code. However, these systems place users in the role of performance engineers, requiring them to make complex decisions about program structure and memory layouts to write programs that run efficiently.
This work presents Galley, a system for declarative tensor programming that allows users to write efficient tensor programs without making complex algorithmic decisions. Galley is the first system to perform cost based lowering of sparse tensor algebra to the imperative language of sparse tensor compilers, and the first to optimize arbitrary operators beyond $\sum$ and *. It does this with a two-level, cost-based program optimizer. At the logical level, it decomposes the input program into a sequence of aggregation steps through a novel extension of the FAQ framework. At the physical level, Galley optimizes and converts each aggregation step to a concrete program, which is compiled and executed with a sparse tensor compiler. We show that Galley produces programs that are 1300×1-300$\times$1 - 300 × faster than competing methods for machine learning over joins and 520×5-20$\times$5 - 20 × faster than a state-of-the-art relational database for subgraph counting workloads. Finally, we show that Galley introduces a minimal optimization overhead across all workloads.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Berlin, Germanyisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Query optimizationccs: Software and its engineering Domain specific languagesccs: Mathematics of computing Mathematical software
@@ 1. Introduction
In recent years, the tensor programming model has become ubiquitous across different fields of computation. Popularized by its use in deep learning, it has been applied to problems such as relational query processing (He et al., 2022; Asada et al., 2022; Koutsoukos et al., 2021), data cleaning (Sun et al., 2023), graph algorithms (Szárnyas et al., 2021), and scientific computing (Wang et al., 2022; Stonebraker et al., 2013; Modi et al., 2021), among others. This model promises a high-level imperative abstraction that is highly efficient as long as the problem can be posed in terms of tensor (i.e., array) operations. It thereby insulates the user from the many low-level concerns that are crucial to a program’s efficient performance. This, in turn, has allowed experts in fields like data science and machine learning (ML) to leverage a wide range of hardware infrastructures without having to become expert in high-performance computing.
Traditional tensor processing frameworks are built on collections of hand-optimized functions, called kernels, which each compute an operation over dense tensors (Abadi et al., 2016; Paszke et al., 2019; Harris et al., 2020; Anderson et al., 1990). The operation can be simple, like a matrix-vector multiplication, or it can be a fusion of multiple semantic operations, like $A(B+C)$. However, most data is fundamentally sparse (i.e., most entries are a fill value like 00), including graph data, one-hot encodings, relational data, 3D physics meshes, sparse neural networks, and others. Even materializing these datasets as dense arrays can be prohibitively costly, so it is crucial to perform the computation over its compressed, sparse format (Kjolstad et al., 2017).
To support sparse data, each framework offers many distinct implementations of each operation, one for each combination of input formats. Each input can either be stored densely or in one of many sparse formats, e.g., CSR, CSC, COO. This has lead to an explosion of required kernels as the number of operations and formats continue to multiply. Understandably, these frameworks have been unable to keep pace with this implementation effort, resulting in spotty coverage for operations over sparse data (Ivanov et al., 2023).
@@
Consider logistic regression over $n$ data points and $d$ features, where $X\in\mathbb{R}^{n\times d}$ is the feature matrix and $\theta\in\mathbb{R}^{d}$ is the parameter vector. Inference is defined as
(1) $P_{i}=\sigma(\sum_{j}X_{ij}\theta_{j})$
where $\sigma$ is the sigmoid function. To compute $Y_{i}=\sum_{j}X_{ij}\theta_{j}$ in the dense case, one could use the ’matmul’ function from the Numpy library. On these inputs, this function specializes to an efficient, hand-coded implementation of dense matrix-vector multiplication. However, if any combination of $X$, $\theta$, and $Y$ could be sparse, Numpy needs to provide eight distinct implementations of matrix-vector multiplication.
To address this issue, the scientific computing community has adopted sparse tensor compilers (STCs) (Kjolstad et al., 2017; Bik et al., 2022; Ahrens et al., 2024; Shaikhha et al., 2022; Hu et al., 2019). These compilers take as input a high-level imperative tensor program and a separate description of the input tensors’ formats. [^1] Some systems separate declarative and imperative concerns with a scheduling language. However, the user still controls both aspects. They automatically produce an efficient kernel implementation in low-level code (LLVM, MLIR, etc.). Thus, STCs offer a form of data independence by letting users separately specify the algorithm and the data layout. For example, Fig. 1 shows a kernel definition for Ex. 1.1 written in Finch, an STC language (Ahrens et al., 2024). The Dense/Sparse input formats are specified in line 1 and are separate from the imperative code in lines 5-15.
However, STCs require users to make a series of complex decisions to produce an efficient implementation. First, users must break their program into a sequence of aggregations. Suppose that the feature matrix in Ex. 1.1 is the result of a matrix multiplication $X=A\cdot B$, so Eq. (1) becomes $P_{i}=\sigma(\sum_{jk}A_{il}B_{lj}\theta_{j})$. Users must choose to sum over $j$ or $l$ first, resulting in vastly different runtimes (Ahrens et al., 2022; Zhang et al., 2021). Then, for each aggregate, they must choose the output format, loop order, and iteration algorithm.
Consider Fig 1. Here, the user must choose the output format for the intermediate R (line 3). In this case, she chose a Dense rather than a Sparse format, which would be 10$\approx$ 10$\times$ 10 × slower. Then, the user chooses the loop order (lines 5-6). In this case, she chose $i$-then-$j$, which is asymptotically faster than $j$-then-$i$ because each out-of-order access to $X$ requires a full scan of the tensor. Finally, the user picks a merge algorithm for each loop that describes how to iterate through the non-zero indices (line 8). Here, $X$ is iterated through, and each non-zero $j$ is looked up in $\theta$. If she chose to iterate through $\theta$, each inner loop would scan the entire vector. Thus, while STCs let them separate the algorithm from the data formats, users are still responsible for optimizing their program to achieve efficient performance.
 0. # Specified format for input tensors
@@ 15. END
In this paper, we propose Galley, a system for declarative sparse tensor programming. Galley makes algorithmic decisions on the users’ behalf, freeing them to focus on the high-level semantics of their program without sacrificing computational efficiency. It accepts input programs written in a declarative language based on Einsum notation, similar to Equation (1), and automatically produces an optimized STC implementation. To do so, it first restructures the program into a sequence of aggregation steps, minimizing total computation and materialization costs (Sec. 5). It then optimizes each step by selecting the loop order, the optimal formats for all intermediate tensors, and the intersection algorithm for each loop (Sec. 6). These decisions are all guided by a system for estimating sparsity via statistics on the input tensors (Sec. 7). Galley builds on fundamental principles from cost-based query optimization while developing new techniques that are specific to producing optimized code for sparse tensor compilers(Khamis et al., 2016).
In the first phase of program optimization, Galley’s logical optimizer rewrites the input program into a sequence of aggregation steps using our new extension of the variable elimination (VE) framework. By adopting this framework, we reduce a complex rewriting problem to one of finding the optimal order to aggregate indices. However, VE has previously been defined only for sum-product expressions, while Galley lets users write arbitrary sparse tensor algebra programs. These programs include non-distributive functions (e.g., $\sum_{i}max(A_{ij},B_{j})$), non-commutative aggregates (e.g., $\sum_{j}\max_{k}B_{jk}$), and other constructs that break the typical assumptions of VE.
Fortunately, we show that this complexity can be neatly managed by reasoning over the annotated expression tree, which lets Galley benefit from algebraic properties when they exist and respect them when they do not. In Sec. 8, we show that these logical optimizations produce up to a $\mathbf{100\times}$ speedup for ML algorithms over composite feature matrices. At the end of this phase, our program has been converted to a sequence of aggregates over pointwise expressions, i.e., the Logical Plan dialect in Fig. 4.
Next, Galley’s physical optimizer generates an efficient STC kernel for each aggregate by choosing the loop order, output format, and merge algorithm. Because sparse outer loops avoid inner iterations, an efficient loop ordering can significantly improve performance. Galley finds the optimal loop order using a novel combination of branch-and-bound and dynamic programming techniques. Following the modern tensor format abstraction (i.e., the fibertree abstraction), Galley then chooses a format for each dimension of the output by examining both the sparsity and write pattern (random vs sequential)(Sze et al., 2017; Chou et al., 2018; Bik et al., 2022). Finally, for each loop index, it selects one input to iterate over and performs lookups on the others.
Galley’s optimizations are guided by a sparsity estimation framework. At its core is a minimal interface for composing statistics. To incorporate a new sparsity estimator in Galley, we only need to implement 5 functions, those for: (1) collecting statistics on inputs, (2-3) handling conjunctive (i.e., *) and disjunctive (i.e., $+$) point-wise operations, (4) handling aggregates, and (5) producing sparsity estimates. We demonstrate this by implementing a uniform sparsity estimator and a worst-case estimator based on cardinality bounds.
In Example 1.1, we assumed that $X$ is given directly. However, feature matrices are generally assembled from more basic inputs. As we show in the following example, Galley can take advantage of this kind of program structure to produce superior performance.
@@
Assume $S_{ipc}\in\mathbb{B}^{n_{i}\times n_{p}\times n_{c}}$ is a boolean sparse tensor, where each non-zero entry is a sale, $i$, of product, $p$, to customer, $c$. $P_{pj}\in\mathbb{R}^{n_{p}\times d}$ and $C_{cj}\in\mathbb{R}^{n_{c}\times d}$ are smaller matrices holding numeric features about products and customers. The feature dimension, indexed by $j$, represents both product and customer features, but each feature comes from either $P_{pj}$ or $C_{cj}$. So, each column is non-zero in either $P$ or $C$, and they are concatenated by addition. $X$ is now defined as $X_{ij}=\sum_{pc}S_{ipc}(P_{pj}+C_{cj})$. Galley’s logical optimizer can take advantage of $X$’s structure by pushing $\theta$ down into its definition, as follows.
$Y_{i}=\sigma(\sum_{jpc}S_{ipc}(P_{pj}\theta_{j}+C_{cj}\theta_{j}))$
We can then push down the summation over $j$, as follows,
(2) $Y_{i}=\sigma(\sum_{pc}S_{ipc}(\sum_{j}P_{pj}\theta_{j}+\sum_{j}C_{cj}\theta_{j}))$
In this way, Galley materializes only vector intermediates, not the full feature matrix. On similar examples, Sec. 8 shows that such changes can yield up to 300×300$\times$300 × faster execution.
After logical optimization, Galley’s physical optimizer produces a single STC kernel for each aggregate. For $\sum_{j}P_{pj}\theta_{j}$, it would produce
    I1 = Dense()
    FOR p=_
        FOR j=_
            I1[p] += P[p::iter,j::iter]*$\theta$[j::lookup]
        END
    END
@@
  • Galley performs cost-based logical optimization with a novel extension of the variable elimination framework to handle arbitrary aggregations and pointwise operators (Sec.5). Galley performs cost-based physical optimization to determine loop orders, tensor formats, and merge algorithms for each dimension (Sec.6).
  • We propose a minimal interface for sparsity estimation to guide optimizations and demonstrate it by implementing two estimators (Sec.7).
  • We evaluate Galley and show that it is 1-300x faster than hand-optimized kernels for mixed dense-sparse workloads and 5-20x faster than a state-of-the-art database for highly sparse workloads based on several evaluation workloads (Sec.8).
2. Related Work
@@ 3.1. Tensor Index Notation
To capture the diverse workloads of tensor programming, we additionally allow the use of arbitrary functions for both aggregates and pointwise operations, nesting aggregates and pointwise operations, and defining multiple outputs. For example, a user could perform logistic regression to predict entities that might be laundering money. Then, they could filter this set based on whether the entities occur in a triangle in the transactions graph. This is represented by $\max_{jk}(E_{ij}E_{jk}E_{ik})$, which is $1$ if $i$ occurs in at least one triangle and 00. This can be written in tensor index notation as:
[…]
@@ C.   Physical Plan
Input := Tns[PIdx...]  Alias := Name[PIdx...]
@@
The input to a Sparse Tensor Compiler is a high-level domain specific language (DSL); it consists of for-loops, in-place aggregates (e.g., +=+=+ =), and arithmetic over indexed tensors (e.g., $A[i,j]*B[j,k]$). Crucially, the for-loops in these expressions are not executed in a dense manner. Instead, these compilers analyze the input formats and the algebraic properties of the expression to determine which index combinations will produce non-fill entries. In Fig. 1, because 00 is the annihilator of multiplication (i.e., $x*0=0$), only the values of $i$ that map to non-zero entries in $X$ and $\theta$ are processed. All other index values will return a zero. So, the outer loop is compiled to an iteration over the intersection of the non-zero $i$ indices in $X$ and $\theta$; Fig. 3 shows how this is simply co-iteration over the top levels of their formats. The inner loop then iterates over the $j$ indices that are non-zero in $X[i,\_]$, i.e., the non-zero columns that occur in each row.
@@ 4. Galley Overview
The input program dialect is equivalent to the tensor index notation defined in Sec. 3.1. Pointwise functions such as $A_{ij}*B_{jk}$ are represented with Map. Aggregates such as $\sum_{i}$ are denoted by Agg. Each assignment is a Query, and previous assignments are referenced with an Alias. Our logistic regression example from Eq. (1) is defined in this dialect as
@@
Note that this notation is compatible with array APIs like Numpy that do not have named indices. Operations like ’matmul’ can be automatically mapped into this language by generating index names for inputs on the fly and renaming whenever operations imply equality between indices. Also, note that aggregates can be over multiple indices, e.g., Agg(+, i, j, k, ...).
@@ 4.1. Logical Plan
The first task in our optimization pipeline, handled by the logical optimizer, breaks down the input program into a sequence of simple aggregates. This is enforced by converting the input program ( 4.A) to a logical plan ( 4.B). This dialect is a restriction of the input dialect, where each query contains a single aggregate statement that wraps an arbitrary combination of Map, Input, and Alias statements. Intuitively, each logical query corresponds to a single STC kernel that produces a single intermediate tensor, but it does not specify details like loop orders and output formats. To perform this conversion soundly, each input query must correspond to a logical query, which produces a semantically equivalent output. To do this efficiently, it must minimize the total cost of all queries in the logical plan.

### v5

@@ Abstract.
The tensor programming abstraction is a foundational paradigm which allows users to write high performance programs via a high-level imperative interface. Recent work on sparse tensor compilers has extended this paradigm to sparse tensors (i.e., tensors where most entries are not explicitly represented). With these systems, users define the semantics of the program and the algorithmic decisions in a concise language that can be compiled to efficient low-level code. However, these systems still require users to make complex decisions about program structure and memory layouts to write efficient programs.
This work presents Galley, a system for declarative tensor programming that allows users to write efficient tensor programs without making complex algorithmic decisions. Galley is the first system to perform cost based lowering of sparse tensor algebra to the imperative language of sparse tensor compilers, and the first to optimize arbitrary operators beyond $\sum$ and *. First, it decomposes the input program into a sequence of aggregation steps through a novel extension of the FAQ framework. Second, Galley optimizes and converts each aggregation step to a concrete program, which is compiled and executed with a sparse tensor compiler. We show that Galley produces programs that are $1-300\times$ faster than competing methods for machine learning over joins and $5-20\times$ faster than a state-of-the-art relational database for subgraph counting workloads with a minimal optimization overhead.
@@
copyright: ccjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 3 (SIGMOD)article: 164publicationmonth: 6doi: 10.1145/3725301ccs: Information systems Query optimizationccs: Software and its engineering Domain specific languagesccs: Mathematics of computing Mathematical software
@@ 1. Introduction
In recent years, the tensor programming (eq. array programming) model has become ubiquitous for high performance computing tasks. It has been applied to problems such as deep learning (Abadi et al., 2016; Paszke et al., 2019; Harris et al., 2020; Anderson et al., 1990), data cleaning (Sun et al., 2023), graph algorithms (Szárnyas et al., 2021), relational query processing (He et al., 2022; Asada et al., 2022; Koutsoukos et al., 2021), and scientific computing (Wang et al., 2022; Stonebraker et al., 2013; Modi et al., 2021), among others. While this approach was originally limited to dense arrays, data from many domains is fundamentally sparse (i.e., most entries are a fill value like 0), including graph data, one-hot encodings, relational data, 3D physics meshes, sparse neural networks, and others. However, modern tensor programming systems like NumPy, PyTorch, and sparse tensor compilers lack the advanced optimization capabilities of relational databases. Instead, users are forced to optimize their programs manually which is challenging and time consuming. In this work, we address this by introducing Galley, a system for declarative sparse tensor programming powered by advanced, cost-based program optimization.
Efficiently processing sparse tensors is challenging. Traditional tensor processing frameworks are collections of hand-optimized functions over dense tensors (Abadi et al., 2016; Paszke et al., 2019; Harris et al., 2020; Anderson et al., 1990). To take advantage of sparsity, these frameworks need to provide implementations for every combination of input tensors’ formats, resulting in spotty coverage for operations over sparse data (Ivanov et al., 2023). Sparse tensor compilers (STCs) have been developed to automatically produce these implementations (Kjolstad et al., 2017; Bik et al., 2022; Ahrens et al., 2024; Shaikhha et al., 2022; Hu et al., 2019). However, these compilers expose even more performance decisions than traditional frameworks, and they similarly lack automatic optimization capabilities. [^1] Some systems separate declarative and imperative concerns with a scheduling language. However, the user still controls both aspects. For a more detailed description of the prototypical STC, we direct the reader to (Kjolstad et al., 2017).
@@
Consider Fig. 1 which implements logistic regression inference in the language of Finch, an STC(Ahrens et al., 2024). Here, the user must choose the output format for the intermediate R (line 3). In this case, she chose a Dense rather than a Sparse format, which would be $\approx 10\times$ slower. Then, the user chooses the loop order (lines 5-6). In this case, she chose $i$-then-$j$, which is asymptotically faster than $j$-then-$i$ because each out-of-order access to $X$ requires a full scan of the tensor. Finally, the user picks a merge algorithm for each loop that describes how to iterate through the non-zero indices (line 8). Here, $X$ is iterated through, and each non-zero $j$ is looked up in $\theta$. If she chose to iterate through $\theta$, each inner loop would scan the entire vector. Even for a simple kernel, these decisions represent a minefield of potential slowdowns.
 0. # Manually specified format for input tensors
@@ 15. END
In this paper, we propose Galley, a system for declarative sparse tensor programming. Galley makes algorithmic decisions on the users’ behalf, freeing them to focus on the high-level semantics of their program without sacrificing computational efficiency. It accepts input programs written in a declarative language, equivalent to the core of the NumPy API, and automatically produces an optimized STC implementation using the Finch compiler (Ahrens et al., 2024). To do so, it first restructures the program into a sequence of aggregation steps, minimizing total computation and materialization costs (Sec. 4). It then optimizes each step by selecting the loop order, the optimal formats for all intermediate tensors, and the merge algorithm for each loop (Sec. 5). These decisions are all guided by a system for estimating sparsity via statistics on the input tensors (Sec. 6). Galley builds on fundamental principles from cost-based query optimization while developing new techniques that are specific to producing optimized code for sparse tensor compilers.
Designing Galley required overcoming three key challenges. First, the high-level optimization requires a complex rewriting of the original program which must respect the algebraic properties of the program. We addressed this by introducing a novel extension of the FAQ framework that can handle arbitrary sparse tensor programs (Khamis et al., 2016). Second, STCs provide a vast design space for kernel implementations which makes the per-aggregate optimization challenging. Galley’s physical optimizer searches through this space efficiently by separating concerns (loop order, output format, and intersection algorithm) and applying branch-and-bound optimization. Lastly, the computational cost of a sparse tensor program depends on the data distribution of the input data which complicates the optimization process. Galley produces these data-dependent cost estimates by leveraging the similarity of sparsity estimation and relational cardinality estimation. By overcoming these challenges, we have attempted to design Galley for a broad set of use cases ranging from sparse ML to graph algorithms and scientific simulations. To this end, we have incorporated Galley into the PyData/Sparse library which implements the full NumPy API for sparse arrays (Abbasi, 2018; Harris et al., 2020).
@@
Let $A$, $B$, and $C$ be sparse matrices, and suppose that you want to compute the matrix chain $ABC$. Because they do not consider the sparsity of the inputs, traditional systems will always perform this in the order $(AB)C$ where the intermediate, $AB$, is stored as a sparse matrix. When given this problem, Galley will optimize at runtime for the input’s sparsities. This allows it to consider plans that are only efficient for specific inputs. For example, it may: 1) re-order the operations to perform $BC$ before multiplying with $A$ 2) store the intermediate as a dense matrix 3) transpose $B$ to iterate over the shared dimension first. In Sec. 7, we show that this can provide a 10x speedup over state-of-the-art tensor frameworks for this example.
@@
  • Galley performs cost-based logical optimization with a novel extension of the variable elimination framework to handle arbitrary aggregations and pointwise operators (Sec.4). Galley performs cost-based physical optimization to determine loop orders, tensor formats, and merge algorithms (Sec.5).
  • We propose a minimal interface for sparsity estimation to guide optimizations and implement two estimators (Sec.6).
  • We evaluate Galley and show that it is 1-300x faster than hand-optimized kernels for mixed dense-sparse workloads and .25-100x faster than a SOTA database for highly sparse workloads (Sec.7).
  • We have implemented Galley as part of the PyData/Sparse sparse array project and the Finch tensor compiler(Pydata, [n. d.]; Ahrens et al., 2024).
@@ 2.1. Tensor Index Notation
To capture the diverse workloads of tensor programming, we additionally allow the use of arbitrary functions for both aggregates and pointwise operations, nesting aggregates and pointwise operations, and defining multiple outputs. For example, a user could perform logistic regression to predict entities that might be laundering money. Then, they could filter this set based on whether the entities occur in a triangle in the transactions graph. This is represented by $\max_{jk}(E_{ij}E_{jk}E_{ik})$, which is $1$ if $i$ occurs in at least one triangle and 0. This can be written in tensor index notation as:
$L_{i}$ $=\sigma(\sum_{j}X_{ij}\theta_{j})>.5$
@@ C.   Physical Plan
Input := Tensor[PIdx...]  Alias := Name[PIdx...]
@@
The input to a Sparse Tensor Compiler is a high-level domain specific language (DSL); it consists of for-loops, in-place aggregates (e.g., +=+=), and arithmetic over indexed tensors (e.g., A[i,j]B[j,k]A[i,j]*B[j,k]). Crucially, the for-loops in these expressions are not executed in a dense manner. Instead, these compilers analyze the input formats and the algebraic properties of the expression to determine which index combinations will produce non-fill entries. In Fig. 1, because 0 is the annihilator of multiplication (i.e., x0=0x*0=0), only the values of $i$ that map to non-zero entries in $X$ and $\theta$ are processed. All other index values will return a zero. So, the outer loop is compiled to an iteration over the intersection of the non-zero $i$ indices in $X$ and $\theta$; Fig. 3 shows how this is simply co-iteration over the top levels of their formats. The inner loop then iterates over the $j$ indices that are non-zero in X[i,_]X[i,\_], i.e., the non-zero columns that occur in each row.
@@ 3. Galley Overview
3.1. Input Program Space
The input program dialect is equivalent to the tensor index notation defined in Sec. 2.1. Pointwise functions such as $A_{ij}*B_{jk}$ are represented with Map. Aggregates such as $\sum_{i}$ are denoted by Agg. Each assignment is a Query, and previous assignments are referenced with an Alias. Crucially, the Op terminal used in both Map and Agg can be any user defined function (e.g. f(x,y)=sin(1+xy)f(x,y)=sin(1+x*y)) as long as it accepts the correct number of arguments (i.e. the number of expressions in the Map and two arguments in Agg). Galley takes advantage of properties of these functions during optimization, specifically distributivity, commutativity, associativity, identity, idempotency, and the existence of an annihilator. Further, users can declare these properties to Galley at runtime. This extensibility is a benefit of Galley’s formal framework. Lastly, Idxs are named symbols (e.g. $i$, $j$), and Tensors are memory-resident input tensors. Our logistic regression example from Fig. 1 is defined in this dialect as
@@ 3.1. Input Program Space
Note that this notation is compatible with array APIs like Numpy that do not have named indices. Operations like ’matmul’ can be automatically mapped into this language by generating index names for inputs on the fly and renaming whenever operations imply equality between indices.
@@ 3.2. Logical Plan
The first task in our optimization pipeline, handled by the logical optimizer, breaks down the input program into a sequence of simple aggregates. This is enforced by converting the input program ( 4.A) to a logical plan ( 4.B). This dialect is a restriction of the input dialect, where each query contains a single aggregate statement that wraps an arbitrary combination of Map, Input, and Alias statements. Intuitively, each logical query corresponds to a single STC kernel that produces a single intermediate tensor, but it does not specify details like loop orders and output formats. To perform this conversion soundly, each input query must correspond to a logical query, which produces a semantically equivalent output. To do this efficiently, Galley must minimize the total cost of all queries in the logical plan.